Please help me again: You draw one card from a standard deck of playing cards. If you draw a diamond, you will win $200.if you draw a face card, which is not diamond, you will win $160. If you draw any othet card,you lose $120. Do you still want to play? Explain
Answer by Boreal(15235)

E(x)=xp(x)
for diamond it is 1/4
face card not a diamond: there are 12 face cards, and nine aren't diamonds.
22 cards are accounted for, all the face cards (12) and the non-face card diamonds (10)
32 cards are left.
E(x)=(1/4)200+(9/52)*160-(32/52)(120)
=50+27.69-73.85=+$3.84.
Whether one wishes to play is more a psychological decision, but if I had enough money I would definitely want to play this, because my expected vale would be nearly $4 each time I did.